# Lie algebra homology with coefficients tensor products of the adjoint representation in relative polynomial degree $2$

### v1

@@ 3. The adjoint action and a reduction
The main purpose of this section is to reduce the calculation of $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ (for $r\geq 2$) to that of the kernel of the morphism $\overline{\delta}_{r}$ introduced in Notation 3.7; this is achieved in Proposition 3.9.
@@ 3.1. The adjoint action
It follows that one can consider the Lie algebra homology: $H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ and this takes values in $\mathfrak{S}_{r}$-modules, naturally with respect to $V$. This is given as the homology of the complex (placed in homological degrees $1$ and 00)
@@
$H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})\cong\left\{\begin{array}[]{ll}\mathrm{coker}(\overline{\mathrm{ad}}_{r})&r=0\\ \ker(\overline{\mathrm{ad}}_{r})&r=1\\ 0&\mbox{otherwise}.\end{array}\right.$
@@
$0\rightarrow\mathbb{C}\mathrm{ycLie}(V)\rightarrow V\otimes\mathbb{L}\mathrm{ie}(V)\stackrel{{\scriptstyle\overline{\mathrm{ad}}_{1}}}{{\rightarrow}}\mathbb{L}\mathrm{ie}(V)\rightarrow V\rightarrow 0$
Here, $\mathbb{C}\mathrm{ycLie}(V)$ is the Schur functor that is associated to the cyclic Lie operad. (Cf. [Lev02], where a stronger result is established, working over $\mathbb{Z}$).
To explain this exact sequence, recall that, if $\mathrm{Lie}(n)$ denotes the $n$th Lie representation (corresponding to the functor $V\mapsto\mathbb{L}\mathrm{ie}(V)^{[n]}$ by the Schur correspondence) and $\mathrm{CycLie}(n)$ the analogous module corresponding to $V\mapsto\mathbb{C}\mathrm{ycLie}(V)^{[n]}$, then these are related by
$\mathrm{Lie}(n)=\mathrm{CycLie}(n+1)\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}$
(this identification explains our conventions for indexing cyclic operads). The collection of morphisms $\mathrm{CycLie}(n+1)\rightarrow\mathrm{Lie}(n)\uparrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}$ obtained by adjunction induce the natural transformation $\mathbb{C}\mathrm{ycLie}(V)\rightarrow V\otimes\mathbb{L}\mathrm{ie}(V)$ above and it is straightforward to check that one obtains a sequence as given. Exactness can be checked at the level of representations of the symmetric groups by using restriction, exploiting the fact that $\mathrm{Lie}(n+1)\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}\cong\mathbb{Q}\mathfrak{S}_{n}$.
@@
The case $n\leq r$ is clear, since $(V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[n]}$ is zero for $n\leq r$. The case $n=r+1$ is shown as follows. The case $r=0$ is immediate. For $r\geq 1$, one uses the the fact that (for $r\geq 1$) $\overline{\mathrm{ad}}_{r}^{[r+1]}$ identifies as an explicit map
@@ 3.2. Reduction of the problem
This can be described explicitly by using the identifications in Example 2.6.
@@
[…]
@@
$\delta^{\prime}_{r}$ ::: $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$
$\delta^{\prime\prime}_{r}$ ::: $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}}.$
Using the $\mathfrak{S}_{r}$-equivariance, these maps can be described by specifying one of their components, as follows.
The morphism underlying $\delta^{\prime}_{r}$ identifies as
$\bigoplus_{i=1}^{r}V\otimes(V^{\otimes i-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-i})\rightarrow\bigoplus_{j=1}^{r}(V^{\otimes j-1}\otimes\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-j}).$
The component with $i=j=1$ is given by
@@
obtained by applying $\overline{\mathrm{ad}}_{1}$ to the first two tensor factors of the domain.
Likewise, the morphism underlying $\delta^{\prime\prime}_{r}$ identifies as
$\bigoplus_{i=1}^{r}V\otimes(V^{\otimes i-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-i})\rightarrow\bigoplus_{1\leq j<k\leq r}(V^{\otimes j-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes k-j-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-k}).$
The component with $i=j=1$ and $k=2$ is
$V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\rightarrow\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2}$
obtained by applying $\overline{\mathrm{ad}}_{1}$ to the first and third tensor factors of the domain.
@@
Notation 3.7.
@@
The underlying functors of the domain and codomain of $\overline{\delta}_{r}$ identify as follows:
@@
For $r\geq 2$, $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ is isomorphic to $\ker\overline{\delta}_{r}$.
@@ 3.3. A geometric visualization
It is useful to visualize the maps in $\overline{\delta}_{r}$ using the $1$-skeleton of an $r-1$-simplex, considered as a directed bipartite graph as follows. The term $V^{\otimes i}\otimes\Lambda^{3}(V)\otimes V^{\otimes r-1-i}$ labels the black vertex of the $(r-1)$-simplex labelled by $i\in\{0,\ldots r-1\}$). For each edge of the $1$-skeleton, one adds a white vertex; on the edge between the black vertices $i\leq j$ this is labelled by the term $V^{\otimes i}\otimes\Lambda^{2}(V)\otimes V^{\otimes j-i}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2-j}$. This white vertex receives directed edges as indicated in the example below for $r=3$:
$\scriptstyle{\Lambda^{1}\otimes\Lambda^{2}\otimes\Lambda^{2}}$$\scriptstyle{\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{2}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}}$.$\scriptstyle{\Lambda^{1}\otimes\Lambda^{1}\otimes\Lambda^{3}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{1}\otimes\Lambda^{2}}$
This allows visualization of ‘restriction’ corresponding to faces. For instance, if one forgets all vertices that have either $\Lambda^{2}(V)$ or $\Lambda^{3}(V)$ in the last factor, one is left with:
@@ 4. Analysing $\overline{\delta}_{r}$
This section first explains how to analyse the kernel of $\overline{\delta}_{r}$ (see Proposition 4.15). This is based on the morphism $\alpha:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{1}\otimes\Lambda^{3}$ that is introduced in Section 4.3 (for the case $r=2$), as well as its analogues $\alpha_{i;j}$ for $r\geq 2$, where $1\leq i<j\leq r$ introduced in Section 4.4.

### v4

@@ 3. The adjoint action and a reduction
The main purpose of this section is to reduce the calculation of $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ (for $r\geq 2$) to that of the kernel of the morphism $\overline{\delta}_{r}$ introduced in Notation 3.8; this is achieved in Proposition 3.10.
@@ 3.1. The adjoint action
It follows that one can consider the Lie algebra homology $H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ and this takes values in $\mathfrak{S}_{r}$-modules, naturally with respect to $V$.
Now, for any $\mathbb{L}\mathrm{ie}(V)$-module $M$, the Lie algebra homology $H_{*}(\mathbb{L}\mathrm{ie}(V);M)$ is the homology of the complex $V\otimes M\rightarrow M$ (placed in homological degrees $1$ and 0) given by the restriction of the module structure. (This follows for example from the fact that $H_{*}(\mathbb{L}\mathrm{ie}(V);M)$ is naturally isomorphic to $\mathrm{Tor}_{*}^{U\mathbb{L}\mathrm{ie}(V)}(\mathbb{Q},M)$ and the universal enveloping algebra $U\mathbb{L}\mathrm{ie}(V)$ is naturally isomorphic to the tensor algebra $T(V)$. One concludes by using the projective resolution of $\mathbb{Q}$ as right $T(V)$-modules provided by $V\otimes T(V)\rightarrow T(V)$, where the map is given by multiplication.)
It follows that $H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ is the homology of the complex
@@
$H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})\cong\left\{\begin{array}[]{ll}\mathrm{coker}(\overline{\mathrm{ad}}_{r})&*=0\\ \ker(\overline{\mathrm{ad}}_{r})&*=1\\ 0&\mbox{otherwise}.\end{array}\right$.
@@
(9) $0\rightarrow\mathbb{C}\mathrm{ycLie}(V)\rightarrow V\otimes\mathbb{L}\mathrm{ie}(V)\stackrel{{\scriptstyle\overline{\mathrm{ad}}_{1}}}{{\rightarrow}}\mathbb{L}\mathrm{ie}(V)\rightarrow V\rightarrow 0.$
Here, $\mathbb{C}\mathrm{ycLie}(V)$ is the Schur functor that is associated to the cyclic Lie operad $\mathrm{CycLie}$. Recall that the cyclic Lie operad is constructed as for the Lie operad, but without a privileged ‘root’ univalent vertex; namely $\mathrm{CycLie}(n)$ is given by the $\mathbb{Q}$-linear span of vertex-oriented, uni-trivalent connected graphs, with univalent vertices labelled bijectively by $\{1,\ldots,n\}$, modulo the AS (antisymmetry) and IHX (analogue of Jacobi) relations.
To explain this exact sequence, recall that, if $\mathrm{Lie}(n)$ denotes the $n$th Lie representation (corresponding to the functor $V\mapsto\mathbb{L}\mathrm{ie}(V)^{[n]}$ by the Schur correspondence, just as $\mathrm{CycLie}(n)$ corresponds to $V\mapsto\mathbb{C}\mathrm{ycLie}(V)^{[n]}$), then one has the identification
$\mathrm{Lie}(n)=\mathrm{CycLie}(n+1)\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}},$
which shows how the Lie operad is obtained by designating a ‘root’ univalent vertex. The collection of morphisms $\mathrm{CycLie}(n+1)\rightarrow\mathrm{Lie}(n)\uparrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}$ obtained by adjunction (using that induction is equivalent to coinduction for finite groups) induce the natural transformation $\mathbb{C}\mathrm{ycLie}(V)\rightarrow V\otimes\mathbb{L}\mathrm{ie}(V)$ above and it is straightforward to check that one obtains a sequence as given.
Exactness can be checked at the level of representations of the symmetric groups by using restriction, exploiting the fact that there is an isomorphism of $\mathfrak{S}_{n}$-modules $\mathrm{Lie}(n+1)\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}\cong\mathbb{Q}\mathfrak{S}_{n}$. Namely, if $n\geq 1$, the homogeneous polynomial degree $n+1$ part of (9) corresponds to the sequence of $\mathbb{Q}\mathfrak{S}_{n+1}$-modules:
(10) $\mathrm{CycLie}(n+1)\rightarrow\mathrm{Lie}(n)\uparrow_{\mathfrak{S}_{n}}^{\mathfrak{S}_{n+1}}\rightarrow\mathrm{Lie}(n+1),$
(the contribution from $V$ is zero). The arrows are respectively injective and surjective, so we seek only to establish exactness in the middle. For this it is sufficient to show exactness after restricting to $\mathbb{Q}\mathfrak{S}_{n}$.
By the Pieri rule (or the Mackey formula), $\big(\mathrm{Lie}(n)\uparrow_{\mathfrak{S}_{n}}^{\mathfrak{S}_{n+1}}\big)\downarrow_{\mathfrak{S}_{n}}^{\mathfrak{S}_{n+1}}$ identifies as $\mathrm{Lie}(n)\oplus(\mathrm{Lie}(n)\downarrow_{\mathfrak{S}_{n-1}}^{\mathfrak{S}_{n}}\big)\uparrow_{\mathfrak{S}_{n-1}}^{\mathfrak{S}_{n}}$, hence as $\mathrm{Lie}(n)\oplus\mathbb{Q}\mathfrak{S}_{n}$. Thus, applying the restriction functor $\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}$ to the sequence (10) gives the sequence of $\mathbb{Q}\mathfrak{S}_{n}$-modules
$\mathrm{Lie}(n)\rightarrow\mathrm{Lie}(n)\oplus\mathbb{Q}\mathfrak{S}_{n}\rightarrow\mathbb{Q}\mathfrak{S}_{n}.$
This is visibly short exact, whence the claimed exactness. (The reader may prefer to rephrase the above simply as a dimension argument.) The exactness in the only non-trivial remaining case, corresponding to homogeneous polynomial degree one, is obvious.
Remark 3.2.
Stronger results (working over $\mathbb{Z}$) than those of Example 3.1 are established in [Lev02, Lev06]. Namely, in [Lev06] the author considers $\mathsf{D}(H)$, the kernel of the bracketing map $H\otimes\mathsf{L}(H)\rightarrow\mathsf{L}(H)$, where $\mathsf{L}(H)$ is the free Lie ring on the finitely-generated free abelian group $H$. He also considers $\mathcal{A}^{t}(H)$, defined similarly to $\mathbb{C}\mathrm{ycLie}(V)$, but working over $\mathbb{Z}$. There is a comparison map $\mathcal{A}^{t}(H)\rightarrow\mathsf{D}(H)$ and the author gives precise information on the failure of this to be an isomorphism. In particular, as a consequence of [Lev06, Theorem 2.5], he recovers the fact that $\mathcal{A}^{t}(H)\otimes\mathbb{Q}\rightarrow\mathsf{D}(H)\otimes\mathbb{Q}$ is an isomorphism. This corresponds to the exactness of (9).
@@
The case $n\leq r$ is clear, since $(V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[n]}$ is zero for $n\leq r$. The case $n=r+1$ is shown as follows. The case $r=0$ is immediate. For $r\geq 1$, one uses the fact that (for $r\geq 1$) $\overline{\mathrm{ad}}_{r}^{[r+1]}$ identifies as an explicit map
@@ 3.2. Reduction of the problem
This can be described explicitly as follows:
@@
Proof.
We use the identifications in Example 2.7, in particular $\mathbb{L}\mathrm{ie}(V)^{[0]}=0$ and $\mathbb{L}\mathrm{ie}(V)^{[1]}=V$.
For the first point, we use the isomorphism $\big(V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r}\big)^{[r+2]}\cong V\otimes(\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+1]}$. Then we have the identification
$(\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+1]}\ \cong\ \bigoplus_{\begin{subarray}{c}(a_{1},\ldots,a_{r})\in\mathbb{N}_{*}^{r}\\ \sum_{i}a_{i}=r+1\end{subarray}}\bigotimes_{i=1}^{r}\mathbb{L}\mathrm{ie}(V)^{[a_{i}]},$
where the sum is over sequences of $r$ positive integers that sum to $r+1$. The only solutions to the latter are when exactly one $a_{i}$ is $2$, all the others being equal to $1$. Since $\mathbb{L}\mathrm{ie}(V)^{[2]}\cong\Lambda^{2}(V)$, taking into account the $\mathfrak{S}_{r}$-action, this gives the stated result.
The second identification is proved similarly; in this case, the sum is over sequences $(0<a_{i}\mid 1\leq i\leq r)$ such that $\sum_{i=1}^{r}a_{i}=r+2$. There are two different types of solution: either exactly one of the $a_{i}$’s is equal to $3$ and the others are all equal to $1$; or exactly two of the $a_{i}$’s are equal to $2$ and the others are all equal to $1$. Since $\mathbb{L}\mathrm{ie}(V)^{[3]}\cong\mathbf{S}_{(2,1)}(V)$, the result follows as above. ∎
@@
$(V\otimes\mathbb{L}\mathrm{ie}(V))^{[2]}\cong V\otimes V$ $\twoheadrightarrow$ $\mathbb{L}\mathrm{ie}(V)^{[2]}=\Lambda^{2}(V)$
$(V\otimes\mathbb{L}\mathrm{ie}(V))^{[3]}\cong V\otimes\Lambda^{2}(V)$ $\twoheadrightarrow$ $\mathbb{L}\mathrm{ie}(V)^{[3]}=\mathbf{S}_{(2,1)}(V),$
@@
$\delta^{\prime}_{r}$ $:$ $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$
$\delta^{\prime\prime}_{r}$ $:$ $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}}.$
By Frobenius reciprocity (i.e., using that $\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$ is left adjoint to $\downarrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$), these are equivalent to $\mathfrak{S}_{r-1}$-equivariant maps:
$\widetilde{\delta^{\prime}_{r}}$ $:$ $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\rightarrow(\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$
$\widetilde{\delta^{\prime\prime}_{r}}$ $:$ $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\rightarrow(\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}},$
using the appropriate restricted $\mathfrak{S}_{r-1}$-module structure on the codomains (not indicated in the notation).
The morphism $\widetilde{\delta^{\prime}_{r}}$ is easily described: it is the composite of
@@
(given by applying $\overline{\mathrm{ad}}_{1}$ to the first two tensor factors of the domain) with the obvious $\mathfrak{S}_{r-1}$-equivariant inclusion $\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1}\subset(\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$.
Likewise, the morphism $\widetilde{\delta^{\prime\prime}_{r}}$ identifies as the composite of
$V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\rightarrow\bigoplus_{1<k\leq r}(\Lambda^{2}(V)\otimes V^{\otimes k-2}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-k})$
(the $k$th component of which is given by applying $\overline{\mathrm{ad}}_{1}$ to the first and (k+1)(k+1)st tensor factors of the domain) with the $\mathfrak{S}_{r-1}$-equivariant inclusion of the codomain into $(\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}}$.
@@
It is convenient to write the expressions without the variable $V$; for example, $V^{\otimes r-1}$ is replaced by the functor $T^{r-1}$.
Notation 3.8.
@@
The underlying functors of the domain and codomain of $\overline{\delta}_{r}$ identify explicitly as follows:
@@
For $r\geq 2$, $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ is naturally isomorphic to $\ker\overline{\delta}_{r}(V)$.
@@ 3.3. A geometric visualization
Henceforth, we use without further comment the identifications $\Lambda^{1}(V)=V=T^{1}(V)$, so that, for $\ell\in\mathbb{N}$, $T^{\ell}$ is naturally isomorphic to $(\Lambda^{1})^{\otimes\ell}$, understood to be $\mathbb{Q}$ for $\ell=0$.
It is useful to visualize the maps in $\overline{\delta}_{r}$ using the $1$-skeleton of an r1r-1-simplex, considered as a directed bipartite graph as follows. The term $T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$ labels the black vertex of the (r1)(r-1)-simplex labelled by $i\in\{1,\ldots r\}$ (using this indexing set for compatibility with that used elsewhere). For each edge of the $1$-skeleton, one adds a white vertex; on the edge between the black vertices $i<j$ this is labelled by the term $T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}$. This white vertex receives directed edges as indicated in the example below for $r=3$:
$\scriptstyle{\Lambda^{1}\otimes\Lambda^{2}\otimes\Lambda^{2}}$$\scriptstyle{\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{2}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{1}\otimes\Lambda^{1}\otimes\Lambda^{3}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{1}\otimes\Lambda^{2}}$
This allows visualization of ‘restriction’ corresponding to faces. For instance, if one forgets all vertices that have either $\Lambda^{2}$ or $\Lambda^{3}$ in the last factor, one is left with:
@@ 4. Analysing $\overline{\delta}_{r}$
This section first explains how to analyse the kernel of $\overline{\delta}_{r}$ (see Proposition 4.17). This is based on the morphism $\alpha:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{1}\otimes\Lambda^{3}$ that is introduced in Section 4.3 (for the case $r=2$), as well as its analogues $\alpha_{i;j}$ (for $r\geq 2$ and $1\leq i<j\leq r$) introduced in Section 4.4.